# The Contiguous Art Gallery Problem is in $\Theta(n\log n)$

### v2

@@
Abstract.
Recently, a natural variant of the Art Gallery problem, known as the Contiguous Art Gallery problem was proposed. Given a simple polygon $P$, the goal is to partition its boundary $\partial P$ into the smallest number of contiguous segments such that each segment is completely visible from some point in $P$. Unlike the classical Art Gallery problem, which is NP-hard, this variant is polynomial-time solvable. At SoCG 2025, three independent works presented algorithms for this problem, each achieving a running time of $O(kn^{5}\log n)$ (or $O(n^{6}\log n)$), where $k$ is the size of an optimal solution. Interestingly, these results were obtained using entirely different approaches, yet all led to roughly the same asymptotic complexity, suggesting that such a running time might be inherent to the problem.
We show that this is not the case. In the realRAM-model, the prevalent model in computational geometry, we present an $O(n\log n)$-time algorithm, achieving an $O(kn^{4})$ factor speed-up over the previous state-of-the-art. We also give a straightforward sorting-based lower bound by reducing from the set intersection problem. We thus show that the Contiguous Art Gallery problem is in $\Theta(n\log n)$.
Acknowledgements.
The authors wish to thank Frank Staals and Jack Spalding-Jamieson for their helpful discussions.
Funding.
This work was supported by the the VILLUM Foundation grant (VIL37507) “Efficient Recomputations for Changeful Problems” and by the Carlsberg Foundation, grant CF24-1929.
@@ 1  Introduction
The Art Gallery problem is a classical problem in computational geometry. Given a simple polygon $P$ with $n$ vertices, the task is to compute the smallest set of guards such that every point $p\in P$ is visible to at least one guard. The problem was first posed by Victor Klee at a meeting of the Mathematical Association of America in 1973 and later formalized by Chvátal [chvatal1975combinatorial], who proved that $\lfloor n/3\rfloor$ guards always suffice and that this combinatorial bound is tight.
@@
Since its introduction, the problem has been extensively studied. O’Rourke and Supowit [Rourke1983hard] showed that the problem is NP-hard if $P$ is allowed to have holes, and Lee and Lin [lee1986VertexNPhard] strengthened this result by proving NP-hardness even when $P$ is a simple polygon without holes. Eidenbenz, Stamm, and Widmayer [EidenbenzStammWidmayer2001] established APX-hardness, and Bonnet and Miltzow [Bonnet2020Parametrized] showed that the problem is W[1]-hard when parametrized by the number of guards. Finally, Abrahamsen, Adamaszek, and Miltzow [abrahamsen2018ETRhard] proved that the problem is $\exists\mathbb{R}$-complete.
Problem variants.
@@
The Art Gallery problem has given rise to an exceptionally large number of problem variants. We briefly review three families of problem variants to illustrate two points: (1) there exists a remarkably large number of problem derivatives that are frequently studied, and (2) almost all of these versions are computationally intractable. When a polynomial-time algorithm exists, it typically has a high-degree polynomial running time.
The first family of variants restricts the structure of the polygon $P$. Krohn and Nilsson [KrohnNilsson2012] showed that the problem remains NP-hard when $P$ is $xy$-monotone. Schuchardt and Hecker [SchuchardtHecker1995] established NP-hardness for orthogonal polygons, and Tomás [Tomas2013ThinOrth] extended this to the case of thin orthogonal polygons. Eidenbenz, Stamm, and Widmayer [EidenbenzStammWidmayer2001] proved NP-hardness even when $P$ is restricted to be a 1.5D terrain (an $x$-monotone polygonal chain that is closed by a base edge).
A second line of work constrains where guards may be placed. Lee and Lin [lee1986VertexNPhard] already proved NP-hardness when guards are restricted to the polygon vertices. Rieck and Scheffer [Rieck2024Dispersive] studied a dispersive variant where guards must maintain a minimum distance from one another, and showed NP-hardness as well. Related to this paradigm is the watchman problem, in which one or more guards move along paths inside the polygon. This problem also admits several sub-variants, the majority of which are NP-hard [ChinNtafos1988, DumitrescuToth2012, Mitchell2013]. Those that can be solved in polynomial time require high-polynomial algorithms, with the best known running times of $O(n^{4})$ [ChinNtafos1991, Tan2001, CarlssonJonssonNilsson1999].
@@
A third family of variants modifies the notion of visibility. Bhattacharya, Kumar, and Bodhayan [BHATTACHARYA2017109] considered edge guards, where a point of $P$ is visible if it can be seen from some point on a guarding edge. Mahdavi, Seddighin, and Ghodsi [MAHDAVI2020163] studied edge guards for rectilinear polygons in which new guarding edges may be introduced in $P$. Biedl and Mehrabi [Biedl2017Mehrabi] analyzed rectilinear visibility, where two points see each other if there exists a rectilinear path between them contained in $P$. All of these variants were shown to be NP-hard. Worman and Keil [Worman2007Decomposition] studied the same rectilinear visibility model but restricted $P$ to be rectilinear also; they obtained an $\tilde{O}(n^{17})$ polynomial-time algorithm. Hoorfar and Bagheri [hoorfar2017guarding] considered rectangle vision, where two points $p,q\in P$ see each other if there exists a rectangle contained in $P$ having $p$ and $q$ as opposite corners. This problem can be solved in $\tilde{O}(n^{17})$ time for simple polygons but becomes NP-hard when holes are allowed [hoorfar2017guarding]. Hillberg, Krohn, and Pahlow [Hillberg2022complexity] showed that Art Gallery remains NP-hard when guards can only see to the right. Vaezi, Roy, and Ghodsi [Vaezi2023Reflective] studied visibility with mirrors that allow guards to see around corners, proving that several of these variants are also NP-hard.
Low-polynomial variants.
@@
Only a few Art Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [Lee1979ptimal] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. Hoorfar and Bagheri [HoorfarBagheri2017HiddenGuards] presented a linear-time algorithm for rectangle vision when $P$ is a histogram. Palios and Tzimas [Palios2013-star-cover] restrict each guard’s visibility region to an orthogonal $r$-star and obtained a near-linear time algorithm. Biedl and Mehrabi [BiedlMehrabi2016_rGuardingThinOrthogonal] combined several restrictions—$P$ must be orthogonal [SchuchardtHecker1995], thin [Tomas2013ThinOrth], and vision rectilinear [Biedl2017Mehrabi]—and achieved a linear-time solution. Chen, Estivill-Castro, and Urrutia [ChenEstivillCastroUrrutia1995] gave a linear-time algorithm for 1.5D terrains when guards can only see to the right.
Contiguous Art Gallery
@@
Among the many known variants, some are motivated by geometric simplicity or mathematical curiosity, while others arise naturally from physical considerations. Laurentini [Laurentini1999GuardingWalls] suggested guarding only the boundary of the polygon, based on the observation that in most art galleries, artworks are displayed on walls. This leads to the question: ‘Can we compute a smallest set of guards such that every point on $\partial P$ is visible to at least one guard?’
@@
Laurentini observed that this problem is NP-hard, and Stade [stade2025interiorBoundaryHard] recently proved that it is $\exists\mathbb{R}$-complete. Thomas Shermer proposed the closely-related Contiguous Art Gallery problem. Here, each guard $g$ can only guard a contiguous segment $[u,v]$ of $\partial P$. The goal is to compute the smallest set of such contiguous guards whose visibility intervals together cover $\partial P$ (see Figure 1). At SoCG 2025, three independent submissions presented realRAM algorithms for this problem, each with comparable high-polynomial running times but based on entirely different methods. Let $k$ denote the size of the optimal solution. Biniaz et al. [biniaz2024contiguous] analyzed the combinatorial structure of contiguous guards. They constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ guards and guarantee that there exists an optimal solution that uses one of the guards from $\mathbb{C}$. They show how to compute $\mathbb{C}$, and how to compute from these an optimal solution n $O(kn^{5}\log n)$ total time.
@@
Merrild, Rysgaard, Schou, and Svenning [merrild2024contiguous] gave an $O(kn^{5}\log n)$-time greedy algorithm. For any point $u\in\partial P$, let $\textnormal{{next}}(u)$ denote the farthest point $v$ along $\partial P$ such that there exists a guard $(g,[u,v])$. Starting from $u$, one can recursively apply next until the resulting set $\{[u,v]\}$ covers $\partial P$. Such a sequence defines a revolution, which they show yields a solution of size at most $k+1$. They compute a revolution in $O(n^{2}\log n)$ time and prove that after $O(kn^{3})$ recursive revolutions, one must obtain an optimal solution, leading to an overall $O(kn^{5}\log n)$-time algorithm.
@@
Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover] took an analytical approach based on the same $\textnormal{{next}}(u)$ function. They parametrize the boundary $\partial P$ by $[1,2n+1)$, where points $i$ and $i+n$ correspond to the same vertex of $P$. Then we can treat next as a function from $[1,n+1)$ to $[1,2n+1)$. They proved that $\partial P$ can be partitioned into $O(n^{3})$ intervals, where next is of constant-complexity in each interval. They extend this to an $O(n^{6}\log n)$-time algorithm.
@@
A subsequent joint version [theEnemy] presented an overview of these three approaches but did not provide a unified algorithm. Thus, there exist three independent approaches that use high polynomial time.
Contribution.
@@
At first sight, Contiguous Art Gallery appears to follow the same pattern as most Art Gallery variants: guarding $\partial P$ in general is hard, and the restricted version admits high-polynomial solutions. The existence of three independent near-$O(kn^{5})$ time solutions may even suggest that this is the intrinsic problem complexity. In this paper, we show that this is not the case as we present an $O(n\log n)$-time algorithm. Our realRAM algorithm uses linear space and achieves an $O(kn^{4})$ factor speed-up over the previous state-of-the-art. [^1] A preliminary version of this paper appeared online with an $\tilde{O}(kn^{2})$-time algorithm. We further prove that this bound is optimal for all comparison-based models of computation in which sorting requires $\Omega(n\log n)$ time, which includes the realRAM. Our result thus provides the first near-linear optimal algorithm for a variant of the Art Gallery problem that is neither artificially over-constrained nor degenerate in structure. Our algorithm combines ideas from all previous approaches, each of which contains a useful combinatorial component that can be exploited for efficiency. Prior works enumerated these components. We replace these brute-force steps with efficient sweep-line algorithms, and combine these insights into a unified efficient algorithm.
Biniaz et al. [biniaz2024contiguous] constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ contiguous guards such that there exists an optimal solution that uses one guard from $\mathbb{C}$. We show that a linear-size set suffices. We compute in $O(n\log n)$ time a linear-size set $X$ of points $u\in\partial P$ such that there exists an optimal solution containing a contiguous guard $(g,[u,v])$. Running the revolution algorithm of Merrild et al. [merrild2024contiguous] for each $x\in X$ would then yield an optimal solution in $O(n^{3}\log n)$ time. However, their method effectively enumerates all interval pairs of a certain arrangement in $P$, leading to the high polynomial factor.
@@
We avoid this enumeration by showing how to compute all values $\textnormal{{next}}(x)$ for $x\in X$ in $O(n\log n)$ total time using a second sweep around $\partial P$. Applying this second sweepline algorithm $k$ times yields an $O(kn\log n)$-time algorithm. To remove the remaining factor of $k$, we compute the revolutions for all $x\in X$ implicitly. To achieve this, we use the algebraic view of the next function introduced by Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover], who showed that next can be represented as a piecewise function with $O(n^{3})$ constant-complexity pieces. Their approach constructs these pieces explicitly in $O(n^{3})$ time by partitioning $\partial P$ into intervals where next is defined by three vertices.
We revisit this construction and show that the number of distinct function pieces is in fact much smaller. This part of the paper is the most technical: we prove that the next function can be represented as a piecewise function with only $O(n)$ constant-complexity pieces, and that this representation can be computed in $O(n\log n)$ total time. Using a lazy segment tree evaluation technique due to Aggarwal et al. [aggarwal1989finding], we can then evaluate all revolutions from $x\in X$ simultaneously and obtain the optimal solution in $O(n\log n)$ time using $O(n)$ space. A particular point of emphasis for our approach is that the running time uses a single logarithmic factor. We show that our algorithms must maintain a convex hull under several types of updates. The best known general-purpose dynamic convex hull structure supports updates in $O(\log^{2}n)$ time. We can get by with an existing first-in-first-out structure to avoid any additional log-factors.
@@ 2  Preliminaries
Parameterizing $\partial P$.
We assume that for any integer $i\in[1,n-1]$, the consecutive vertices $v_{i}$ and $v_{i+1}$ of $P$ appear counter-clockwise. Consequently, the interior of $P$ lies immediately left of $\overline{v_{i}\,v_{i+1}}$. For convenience, we define a continuous surjective function $1,2n+1)\rightarrow\partial P$ such that for every vertex $v_{j}$, both $j$ and $n+j$ map to $v_{j}$. Thus, each point on $\partial P$ can be represented by two real values: one in $[1,n+1)$ and one in $[n+1,2n+1)$.
Using this parametrization, we define (open) chains as follows. A chain is a sequence of edges. For two points $u,v\in\partial P$, we denote by $[u,v]$ the chain obtained by traversing $\partial P$ from $u$ to $v$ in counter-clockwise order. We denote by $(u,v)$ the open chain, consisting of all points $x\in[u,v]$ with $x\neq u$ and $x\neq v$; that is, $(u,v):=\{\,x\in[u,v]\mid x\neq u,\,x\neq v\,\}.$ Observe that the chain between any two points on $\partial P$ can be described as $[u,v]$ with $u\in[1,n+1)$ and $v\in[u,2n+1)$.
Problem statement.
A point $x$ sees a point $y$ if the segment $\overline{x\,y}$ is contained in $P$. A (contiguous) guard is a tuple $(g,[u,v])$ consisting of a point $g\in P$ and a chain $u,v]\subset\partial P$ such that every point on the chain is visible from $g$.
@@
We say that a guard $(g,[u,v])$ dominates another guard $(g^{\prime},[u^{\prime},v^{\prime}])$ if $u^{\prime},v^{\prime}]\subset[u,v]$. It strictly dominates $(g^{\prime},[u^{\prime},v^{\prime}])$ if $u^{\prime},v^{\prime}]\subsetneq[u,v]$. For a fixed $u\in[1,n+1)$, we frequently compute a maximal $v\in[u,2n+1)$ such that there exists a point $g\in P$ for which $(g,[u,v])$ is a guard. By this we mean that for all other guards $(g^{\prime},[u,v^{\prime}])$, it holds that $v^{\prime}\in[u,v]$.
Visibility core.
For any chain $[u,v]$ between two points $u,v\in\partial P$, we define the visibility core $\mathcal{E}[u,v]$ as the set of points $p\in\mathbb{R}^{2}$ that lie left of all edges in $[u,v]$. We can use the visibility core to decide whether a single guard suffices.
Observation 1.

### v4

@@
Abstract
Recently, a natural variant of the Art Gallery problem, known as the Contiguous Art Gallery problem was proposed. Given a simple polygon $P$, the goal is to partition its boundary $\partial P$ into the smallest number of contiguous segments such that each segment is completely visible from some point in $P$. Unlike the classical Art Gallery problem, which is NP-hard, this variant is polynomial-time solvable. At SoCG 2025, three independent works presented algorithms for this problem, each achieving a running time of $O(kn^{5}\log n)$ (or $O(n^{6}\log n)$), where $k$ is the size of an optimal solution. Interestingly, these results were obtained using entirely different approaches, yet all led to roughly the same asymptotic complexity.
In the realRAM-model, the prevalent model in computational geometry, we present an $O(n\log n)$-time algorithm, achieving an $O(kn^{4})$ factor speed-up over the previous state-of-the-art. We also give a straightforward lower bound reduction from the set intersection problem. We thus show that the Contiguous Art Gallery problem is in $\Theta(n\log n)$.
keywords:
art gallery problem, optimal algorithm, simple polygon
@@ 1 Introduction
The Art Gallery problem is a classical problem in computational geometry. Given a simple polygon $P$ with $n$ vertices, the task is to compute the smallest set of guards such that every point $p\in P$ is visible to at least one guard. The problem was first posed by Victor Klee in 1973 and later formalized by Chvátal [chvatal1975combinatorial], who proved that $\lfloor n/3\rfloor$ guards always suffice and that this combinatorial bound is tight. Since its introduction, the problem has been extensively studied. O’Rourke and Supowit [Rourke1983hard] showed that the problem is NP-hard if $P$ is allowed to have holes, and Lee and Lin [lee1986VertexNPhard] strengthened this result by proving NP-hardness even when $P$ is a simple polygon without holes. Eidenbenz, Stamm, and Widmayer [EidenbenzStammWidmayer2001] established APX-hardness, and Bonnet and Miltzow [Bonnet2020Parametrized] showed that the problem is W[1]-hard when parametrized by the number of guards. Finally, Abrahamsen, Adamaszek, and Miltzow [abrahamsen2018ETRhard] proved that the problem is $\exists\mathbb{R}$-complete.
@@
Problem variants.  The Art Gallery problem has given rise to an exceptionally large number of problem variants. We briefly review three families of problem variants to illustrate two points: (1) there exists a remarkably large number of problem derivatives that are frequently studied, and (2) almost all of these are computationally intractable. When a polynomial-time algorithm exists, it typically has a high-degree.
@@
The first family of variants restricts the structure of the polygon $P$. Krohn and Nilsson [KrohnNilsson2012] showed that the problem remains NP-hard when $P$ is $x$-monotone. Schuchardt and Hecker [SchuchardtHecker1995] established NP-hardness for orthogonal polygons, and Tomás [Tomas2013ThinOrth] extended this to the case of thin orthogonal polygons. King and Krohn [15D_terrain] proved NP-hardness even when $P$ is restricted to be a 1.5D terrain. A second line of work constrains guard placement. Lee and Lin [lee1986VertexNPhard] proved NP-hardness when guards are restricted to the polygon vertices. Rieck and Scheffer [Rieck2024Dispersive] studied a dispersive variant where guards must maintain a minimum distance from one another, and showed NP-hardness as well. Related is the watchman problem, in which one or more guards move along paths inside the polygon. This problem also admits several sub-variants, the majority of which are NP-hard [ChinNtafos1988, DumitrescuToth2012, Mitchell2013]. A third family of variants modifies the notion of visibility. Lee and Lin [lee1986VertexNPhard] considered edge guards, where a point of $P$ is visible if it can be seen from some point on a guarding edge. Mahdavi, Seddighin, and Ghodsi [MAHDAVI2020163] studied edge guards with orthogonal vision that can see through up to $k$ walls for rectilinear polygons in which new guarding edges may be introduced in $P$. Biedl and Mehrabi [Biedl2017Mehrabi] analyzed rectilinear visibility, where two points see each other if there exists a rectilinear path between them contained in the orthogonal polygon $P$ that may contains holes. All of these variants were shown to be NP-hard. Motwani, Raghunathan, and Saran [s-covering] showed that this problem with rectilinear vision can be solved in $O(n^{8})$ time in a rectilinear polygon without holes. Worman and Keil [Worman2007Decomposition] considered rectangle vision, where two points $p,q\in P$ see each other if there exists a rectangle contained in $P$ having $p$ and $q$ as opposite corners; they obtained an $\tilde{O}(n^{17})$ polynomial-time algorithm.
@@
Low-polynomial variants.  Only a few Art Gallery variants are solvable in low-polynomial time, and all known results rely on very strong structural restrictions. Lee and Preparata [Lee1979ptimal] gave a linear-time algorithm to decide whether a polygon can be guarded by a single guard. De Berg, Durocher, and Mehrabi. [DBLP:journals/jda/BergDM17] presented a linear-time algorithm to minimize the number of edge guards under orthogonal vision when $P$ an rectilinear polygon. Palios and Tzimas [Palios2013-star-cover] restrict each guard’s visibility region to an orthogonal $r$-star and restrict the polygon to class-3 orthogonal polygons, and obtained near-linear time. Biedl and Mehrabi [BiedlMehrabi2016_rGuardingThinOrthogonal] combined several restrictions—$P$ must be orthogonal [SchuchardtHecker1995], thin [Tomas2013ThinOrth], and vision rectilinear [Biedl2017Mehrabi].
@@
Contiguous Art Gallery.  Among the many known variants, some are motivated by geometric simplicity or mathematical curiosity, while others arise naturally from physical considerations. Laurentini [Laurentini1999GuardingWalls] suggested guarding only the boundary of the polygon, based on the observation that in most galleries, artworks are displayed on walls. This leads to the question: ‘Can we compute a smallest set of guards such that every point on $\partial P$ is visible to at least one guard?’
@@
Laurentini observed that this problem is NP-hard, and Stade [stade2025interiorBoundaryHard] recently proved that it is $\exists\mathbb{R}$-complete. Thomas Shermer proposed the closely-related Contiguous Art Gallery problem. Here, each guard $g$ can only guard a contiguous segment $[u,v]$ of $\partial P$. The goal is to compute the smallest set of contiguous guards which together see $\partial P$ (see Figure 1).
At SoCG 2025, three independent submissions presented realRAM solutions, each with comparable high-polynomial running times but based on entirely different methods. Let $k$ denote the size of the optimal solution. Biniaz et al. [biniaz2024contiguous] constructed a candidate set $\mathbb{C}$ of $O(n^{4})$ guards and guarantee that there exists an optimal solution using a guard from $\mathbb{C}$. They compute $\mathbb{C}$, and compute from $\mathbb{C}$ an optimal solution, in $O(kn^{5}\log n)$ total time.
@@
Merrild, Rysgaard, Schou, and Svenning [merrild2024contiguous] gave an $O(kn^{5}\log n)$-time greedy algorithm. For any point $u\in\partial P$, let $\textnormal{{next}}(u)$ denote the farthest point $v$ along $\partial P$ such that there exists a guard that sees $[u,v]$. Starting from $u$, recursively apply next until the resulting set $\{[u,v]\}$ covers $\partial P$. Such a sequence defines a revolution, which they show yields a solution of size at most $k+1$. They compute a revolution in $O(n^{2}\log n)$ time and prove that after $O(kn^{3})$ revolutions, one obtains an optimal solution, leading to an $O(kn^{5}\log n)$-time algorithm.
@@
Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover] took an analytical approach based on the same $\textnormal{{next}}(u)$ function. They parametrize the boundary $\partial P$ by $[1,2n+1)$, where points $i$ and $i+n$ correspond to the same vertex of $P$. Then we can treat next as a function from $[1,n+1)$ to $[1,2n+1)$. $\partial P$ can be partitioned into $O(n^{3})$ intervals, where next is of constant-complexity in each interval. They extend this to an $O(n^{6}\log n)$-time algorithm.
@@
Their joint version [theEnemy] presented an overview of these three approaches but did not provide a unified algorithm. Thus, there exist three approaches that use high polynomial time.
@@
Contribution.  At first sight, Contiguous Art Gallery appears to follow the same pattern as most Art Gallery variants: guarding $\partial P$ in general is hard, and the restricted version admits high-polynomial solutions. Yet, we show that we can significantly improve the running time to $O(n\log n)$. Our realRAM algorithm uses linear space and achieves an $O(kn^{4})$ factor speed-up over the previous state-of-the-art (we note that preliminary version of this paper appeared online with an $\tilde{O}(kn^{2})$-time algorithm [deberg2025simpler]). We further prove that this bound is optimal for all comparison-based models of computation in which sorting requires $\Omega(n\log n)$ time, which includes the realRAM. Our result thus provides a significant speedup and tight bounds for the Contiguous Art Gallery problem.
@@
Technical contribution.  First, we explore the extent that previous techniques can solve this problem. Biniaz et al. [biniaz2024contiguous] constructed a candidate set $\mathbb{C}$ of size $O(n^{4})$ contiguous guards where at least one optimal solution uses at least one guard from $\mathbb{C}$. We show that a linear-size set suffices. We compute in $O(n\log n)$ time a linear-size set $X$ of points $u\in\partial P$ such that there exists an optimal solution containing a contiguous guard $(g,[u,v])$. Running the revolution algorithm of Merrild et al. [merrild2024contiguous] for each $x\in X$ would then yield an optimal solution in $O(n^{3}\log n)$ time. We avoid this brute force by showing how to compute all values $\textnormal{{next}}(x)$ for $x\in X$ in $O(n\log n)$ total time using a second sweep around $\partial P$. Applying this second sweepline algorithm $k$ times yields an $O(kn\log n)$-time algorithm. To remove the remaining factor of $k$, we compute the revolutions for all $x\in X$ implicitly. To achieve this, we use the algebraic view of the next function introduced by Robson, Spalding-Jamieson, and Zheng [RobsonSpaldingZheng2024_AnalyticArcCover], who showed that next can be represented as a piecewise function with $O(n^{3})$ constant-complexity pieces. Their approach constructs these pieces explicitly in $O(n^{3})$ time by partitioning $\partial P$ into intervals where next is defined by three vertices. We show that the number of distinct function pieces is in fact much smaller. This part of the paper is the most technical: we prove that the next function can be represented as a piecewise function with only $O(n)$ constant-complexity pieces, and that this representation can be computed in $O(n\log n)$ total time. This yields an $O(kn+n\log n)$-time algorithm where for each of the $O(n)$ starting guards $x\in X$, we explicitly compute a full revolution in $O(k)$ time by evaluating the function next at most $k+1$ times. We therefore argue that with our $O(n)$ starting guards, and optimal usage of all three techniques, one finds an $O(kn+n\log n)$-time algorithm. We shave the factor $k$ by applying a simultaneous-evaluation technique by Aggarwal et al. [aggarwal1989finding]. This computes for all $x\in X$ their revolution in $O(n\log n)$ total time.
@@ 2 Preliminaries
Parameterizing $\partial P$.  We assume that for any integer $i\in[1,n-1]$, the consecutive vertices $v_{i}$ and $v_{i+1}$ of $P$ appear counter-clockwise. Consequently, the interior of $P$ lies immediately left of $\overline{v_{i}\,v_{i+1}}$. For convenience, we define a continuous surjective function $1,2n+1)\rightarrow\partial P$ such that for every vertex $v_{j}$, both $j$ and $n+j$ map to $v_{j}$. Thus, each point on $\partial P$ can be represented by two real values: one in $[1,n+1)$ and one in $[n+1,2n+1)$.
Using this parametrization, we define (open) chains as follows. A chain is a sequence of edges. For two points $u,v\in\partial P$, we denote by $[u,v]$ the chain obtained by traversing $\partial P$ from $u$ to $v$ in counter-clockwise order. We denote by $(u,v)$ the open chain, consisting of all points $x\in[u,v]$ with $x\neq u$ and $x\neq v$. Observe that the chain between any two points on $\partial P$ can be described as $[u,v]$ with $u\in[1,n+1)$ and $v\in[u,2n+1)$.
Problem statement.  A point $x$ sees a point $y$ if the segment $\overline{x\,y}$ is contained in $P$. A (contiguous) guard is a tuple $(g,[u,v])$ consisting of a point $g\in P$ and a chain $u,v]\subseteq\partial P$ such that every point on the chain is visible from $g$.
@@
We say that a guard $(g,[u,v])$ dominates another guard $(g^{\prime},[u^{\prime},v^{\prime}])$ if $u^{\prime},v^{\prime}]\subseteq[u,v]$. It strictly dominates $(g^{\prime},[u^{\prime},v^{\prime}])$ if $u^{\prime},v^{\prime}]\subsetneq[u,v]$. For a fixed $u\in[1,n+1)$, we frequently compute a maximal $v\in[u,2n+1)$ such that there exists a point $g\in P$ for which $(g,[u,v])$ is a guard. By this we mean that for all other guards $(g^{\prime},[u,v^{\prime}])$, it holds that $v^{\prime}\in[u,v]$.
Visibility core.  For any chain $[u,v]$ between two points $u,v\in\partial P$, we define the visibility core $\mathcal{E}[u,v]$ as the set of points $p\in\mathbb{R}^{2}$ that lie left of all edges in $[u,v]$. We can use this to decide whether a single guard suffices, and henceforth we thus assume the optimal number of guards $k$ to be at least two:
{observation}